Beth's Bakery can comfortably produce 60 brownies in one day. If Beth takes some unusual measures, such as hiring her two aunts to help in the kitchen and work overtime, she can produce up to 100 brownies in one day.

(a) What are the design and effective capacities for Beth's Bakery?

(b) If Beth is currently producing 64 brownies, compute the capacity utilization for both measures. What can you conclude?
Answer #1

a. Effective capacity = 60 brownies

Design capacity = 100 brownies.

b. Utilization effective = actual output/ effective output *100

                                     = 64/60 *100 = 106.67%

Utilization design = actual output / design capacity *100

                              = 64 / 100 * 100 = 64%

the utilization rate show that the facility current outputs is far below its design capacity and higher than its effective capacity. the design capacity occur only with extra help.